Calculate the percentages of elements contained in two specified chemical compounds:

1. a carboxyl group (-COOH)
2. NH3
1
Expert's answer
2020-05-24T03:01:07-0400

1. 12 x 100 / 45= 26,6666667 % of C.

32 x 100 /45= 71,1111111 % of O.

1 x 100/45 = 2,2222222% of H.

2. 14 x 100/17 = 82% of N.

100 - 82 = 18 % of H.